# Extracting and Transferring Abilities For Building Multi-lingual Ability-enhanced Large Language Models

### v1

@@ Abstract
Multi-lingual ability transfer has become increasingly important for the broad application of large language models (LLMs). Existing work highly relies on training with the multi-lingual ability-related data, which may be not available for low-resource languages. To solve it, we propose a Multi-lingual Ability Extraction and Transfer approach, named as MAET. Our key idea is to decompose and extract language-agnostic ability-related weights from LLMs, and transfer them across different languages by simple addition and subtraction operations without training. Specially, our MAET consists of the extraction and transfer stages. In the extraction stage, we firstly locate key neurons that are highly related to specific abilities, and then employ them to extract the transferable ability-specific weights. In the transfer stage, we further select the ability-related parameter tensors, and design the merging strategy based on the linguistic and ability specific weights, to build the multi-lingual ability-enhanced LLM. To demonstrate the effectiveness of our proposed approach, we conduct extensive experiments on mathematical and scientific tasks in both high-resource lingual and low-resource lingual scenarios. Experiment results have shown that MAET can effectively and efficiently extract and transfer the advanced abilities, and outperform training-based baselines methods. Our code and data are available at https://github.com/RUCAIBox/MAET.
@@ 1 Introduction
Large language models (LLMs) have recently shown remarkable performance on various general tasks, e.g., text generation and question answering (Zhao et al., 2023; OpenAI, 2023; Dubey et al., 2024). Despite the success, LLMs are still struggling to solve complex tasks (e.g., mathematical reasoning), which require LLMs to possess specific advanced abilities (e.g., deductive reasoning) and knowledge (e.g., mathematical theory) (Yue et al., 2024; Lu et al., 2022). To address it and further improve LLMs, existing work either collects the related data to train LLMs (Du et al., 2024; Chen et al., 2024a), or merges the parameters of existing well-performed LLMs to transfer their advanced abilities into one single model (Ilharco et al., 2023; Yadav et al., 2023; Yu et al., 2024a).
Despite the success, it is not easy to collect sufficient training corpus or well-trained LLMs related to specific abilities, especially in multi-lingual scenarios. Especially, some popular languages (e.g., English) have dominated the linguistic expressions of the open web data, and the amount of available data for low-resource languages (e.g., Bengali or Telugu) is highly limited (Patzelt, 2024; Mirashi et al., 2024). Fortunately, existing work (Zhao et al., 2024; Schäfer et al., 2024) has revealed that the learned knowledge from one language by LLMs could be inherited and leveraged by other languages. For example, Llama-series LLMs are trained mainly on English texts, while they can also solve the tasks based on other languages. Such a finding has been widely explored in either improving the overall performance of multi-lingual LLMs (Schäfer et al., 2024) or enhancing fine-grained knowledge (Chen et al., 2024a). However, the related work mostly relies on training with ability-related corpus in the target language, which is not always available for low-resource languages.
@@
To assess the effectiveness of our approach, we conduct the evaluation based on two comprehensive reasoning benchmarks, namely Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) as the evaluation benchmarks. According to the evaluation results, the proposed approach MAET outperforms other competitive baseline methods (e.g., continual pre-training (Gururangan et al., 2020) and model merging methods with task vectors (Ilharco et al., 2023), achieving the 9.1% relative improvement upon the backbone LLM. In addition, our approach can work well with relatively fewer training data, demonstrating an improved efficiency in practice. In conclusion, our contribution can be summarized as follows,
(1) Our research has found that advanced abilities can be extracted from the single-lingual corpus and transferred across languages without the multi-lingual ability-related corpus, enabling to efficiently empower LLMs with special advanced abilities.
(2) We propose an effective and efficient approach named MAET, which first identifies and extracts the ability-related weights in LLMs and then only leverages simple addition and subtraction operations to build a multi-lingual ability-enhanced LLM.
(3) Our approach MAET achieves better performance than the competitive baseline methods (e.g., continual pre-training and model merging with task vector) in multi-lingual complex reasoning tasks, including mathematical reasoning tasks and scientific reasoning tasks.
@@ 2 Related Work
We introduce the related work from the following three perspectives:
@@
Continual Pre-training. Although LLMs have shown remarkable performance on various downstream work, they still struggle in several specific tasks, e.g., complex reasoning tasks (Paster et al., 2024; Shao et al., 2024) or low-resource lingual scenarios (Hedderich et al., 2021; Panchbhai & Pankanti, 2021). To adapt LLMs pre-trained on the general corpus to multi-lingual scenarios or specific tasks, existing work (Luo et al., 2022; Taylor et al., 2022; Zhao et al., 2022; Zhang et al., 2024a) has collected the corresponding corpus to continually pre-train (CPT) LLMs. During the continual pre-training process, the mixture strategy between the general corpus and the ability-related corpus should be carefully considered to prevent hurting the general abilities of LLMs (Ye et al., 2024; Xie et al., 2023; Chen et al., 2024a; Siriwardhana et al., 2024). However, previous study Chang et al. (2024); Lu et al. (2023) has pointed out that it is difficult to collect the required corpus, especially for low-resource language corpus. Therefore, synthesizing data from powerful LLMs is widely utilized to expand the task-specific training corpus (Chen et al., 2021b; Yu et al., 2024b; Zhou et al., 2024a). Besides, because of the limitation of computation resources, a series of approaches (Hu et al., 2022; Li & Liang, 2021; Dettmers et al., 2023) only train several parameters to reduce the expenses. In this work, we focus on adapting LLMs to multilingual complex reasoning scenarios through continually pre-training LLMs on the single-lingual task-specific corpus.
@@
Knowledge Editing. According to lottery ticket hypothesis (Frankle & Carbin, 2019), training a small number of model parameters will achieve comparable or even better performance on downstream tasks. Existing study (Du et al., 2024; Wang et al., 2024b; Gong et al., 2024) has leveraged the inner information of LLMs, e.g., gradient or cosine similarity between different hidden states, to select and train the related sub-network. Besides, the probe (i.e., a newly initialized parameter) can be implemented to detect the knowledge of LLMs and process targeted repair (Wang et al., 2024a; Jiang et al., 2024). However, Since these approaches need to calculate and select a sub-network of each training instance, which might cause the instability of the training process, several study (Chen et al., 2024b; Zhang et al., 2024b) pointed out that the task-related sub-network can be determined before the training process, and only updating the value of the corresponding neurons can achieve better performance. In this work, we focus on editing the task-specific neurons of LLMs to improve the corresponding capacities in multi-lingual scenarios.
@@
Model Merging. Given that the CPT process will bring huge computational expenses, previous work leveraged model merging techniques to integrate different abilities (e.g., mathematical reasoning and code synthesizing) into one model (Yang et al., 2024; Xu et al., 2024; Stoica et al., 2024). During the merging process, the interference between different LLMs might be conflict with each other and affect the final performance. Therefore, researchers proposed the clip (Yadav et al., 2023) or randomly dropout (Yu et al., 2024a) mechanism to mitigate the performance decrease. Moreover, the selection of the hyper-parameters (e.g., weight of each model) is the challenge of the model merging process, and previous work (Zhou et al., 2024b; Matena & Raffel, 2022) utilized the inner parameters of LLMs or external matrixes to determine the hyper-parameters. Inspired by the above work, we try to locate the task-related sub-networks of LLMs and transfer the advanced abilities.
@@ 3 Preliminary
Despite that LLMs exhibit excellent performance on general tasks, they still have limited advanced abilities, e.g., mathematical and scientific reasoning abilities. A typical approach to enhancing these abilities is to continually pre-train (CPT) LLMs with ability-related corpus. However, such training data might not always be available or sufficient, especially for minor domains (e.g., Bengali). In this work, we focus on the task of ability extraction and transfer by continual pre-training and merging LLMs. Concretely, LLMs are trained on collected corpus from a certain domain, and we aim to only transfer its learned advanced capabilities to another target domain (Zhuang et al., 2021; Farahani et al., 2021) without further training efforts. In this work, we mainly study the cross-lingual scene where the linguistic-agnostic advanced ability and linguistic abilities should be extracted and transferred, to help build a unified multi-lingual ability-enhanced LLM.
@@ 4 Approach
Figure 1: The framework of our approach MAET, including extraction and transfer stages. In the extraction stage, we first locate the key neurons, and utilize the single-lingual ability-related corpus and general corpus to train the LLM on these neurons to obtain the ability-specific weight. Then, we remove the parameter tensors related to language knowledge in the ability weight and transfer the remaining to the base LLM. After these stages, we can obtain a powerful LLM with advanced abilities that can solve the corresponding tasks in multi-lingual scenarios.
In this section, we propose the Multi-lingual Ability Extraction and Transfer approach, named as MAET, which can effectively transfer the advanced abilities from single-lingual LLMs, to build a multi-lingual ability-enhanced LLM. The key motivation of our approach is to identify and extract ability-related neurons or weights, and transfer the target abilities into a LLM in an efficient way. The overall framework of MAET is presented in Figure 1.
@@ 4.1 Ability-related Weights Extraction
Locating the Key Neurons. The gradient of each neuron in LLMs can be utilized to estimate its correlation degree with specific task ability , we select those with high gradient values as key neurons. To this end, we first use a small amount of ability-related corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ to continually pre-train the LLM. During training, the LLM learns the language modeling task and each neuron is updated by the gradients associated by the training instances. Due to the high cost of calculating the accumulation of gradient at each training step, we calculate the value changes of the LLM neurons before and after the training process to approximate the importance. Formally, the importance function $I(A_{i},\theta_{j})$ of neurons can be computed as:
$I(A_{i},\theta_{j})=\sum_{d_{k}\in\mathcal{C}_{L_{0},A_{i}}^{\prime}}{\text{Gradient}\left(\theta_{j},d_{k}\right)}\approx\lambda\cdot\parallel\tilde{\theta}_{j}-\theta_{j}\parallel,$ (1)
where $d_{k}$ denotes the $k$-th instance of training corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ and $\tilde{\theta}_{j}$ denote the value of the $j$-th neuron of LLM after training, respectively. Based on it, we rank all neurons according to their importance scores, and then select the top $k_{1}\%$ ones into the set $\mathcal{N}_{A_{i}}$ as the key neurons.
Learning Ability-specific Weights. Based on the identified key neurons in $\mathcal{N}_{A_{i}}$, we further learn the ability-specific parameter weights. Our motivation is to decompose the parameter weights according to their changes before and after the LLM has mastered a specific ability, which is achievable owing to the modularity and composition nature of the LLM parameter matrices (Yu et al., 2024a; Shazeer et al., 2017). First, we utilize the key neurons locating method mentioned above to extract the ability-related neuron set $\mathcal{N}_{A_{i}}$, and also obtain the language-related neuron set $\mathcal{N}_{L_{0}}$ via the same way. Then, we train the LLM with the mixture of ability-related corpus and general corpus on the key neuron set $\mathcal{N}_{A_{i}}\bigcup\mathcal{N}_{L_{0}}$ and $\mathcal{N}_{L_{0}}$ respectively, to obtain two specific models, denoted as $\text{LLM}_{A_{i},L_{0}}$ with parameters $\Theta_{A_{i},L_{0}}$ and $\text{LLM}_{L_{0}}$ with parameters $\Theta_{L_{0}}$. Next, we measure the parameter changes between the backbone and the trained models, and obtain the ability-specific weights via the parameter decomposition operation as:
@@
where $\alpha$ and $\beta$ are tunable coefficients to balance the two parts of weight differences, and $\Theta_{o}$ denote the original parameters of the LLM, which serves as the reference for parameter decomposition. As we only train the parameters within the neuron set, its weight difference should preserve the knowledge about the corresponding ability. Thus, it can be regarded as the ability-specific parameter representations, and is promising to transfer the ability into other LLMs by the addition operation.
1
Input : Single-lingual ability-related corpus $\mathcal{C}_{L_{0},A_{i}}$, multi-lingual general corpus $\mathcal{C}_{L_{0}},\mathcal{C}_{L_{1}},\dots,\mathcal{C}_{L_{n}}$, and the parameters of the backbone model $\Theta_{o}$.
Output : A well-trained multi-lingual ability-enhanced LLM.
2
// Ability-related Weights Extraction
3 $\theta^{\prime}\leftarrow\text{CPT}(\mathcal{C}_{L_{0},A_{i}},\Theta_{o}$);
4
5for $j$-th neuron in $\Theta_{o}$ do
6       Calculate the importance score of the corresponding neuron using Eq. 1;
7
8
9Identify the key neuron set $\mathcal{N}_{A_{i}}$;
10
[…]
12
[…]
14
15Learning the ability-specific weight $R(A_{i})$ using Eq. 2;
16
// Multi-lingual Ability Transfer
17
18Obtaining the multi-lingual weight $R_{Lang}$ using Eq. 3;
19
20for $j$-th parameter tensor in LLM do
21       Calculate the correlation using Eq. 4;
22
23
24Identify the ability-related parameters $\mathcal{T}$;
25
26Transfer the ability to multi-lingual scenarios using Eq. 5;
27
28Obtain the well-trained multi-lingual ability-enhanced LLM.
Algorithm 1 The complete procedure of our proposed approach.
@@ 4.2 Multi-lingual Ability Transfer
After obtaining the ability-related weights, in this part, we utilize them to transfer and integrate the abilities, to build a multi-lingual ability-enhanced LLM.
@@
where we use the dot product to implement the similarity function $\texttt{sim}\left(\cdot\right)$. After obtaining the similarity of all tensors, we rank them in a descending order based on the similarity values, and then select the last $k_{2}\%$ parameters into the set $\mathcal{T}$ as the ability-related parameters.
@@
Aspects CPT MoE LoRA MoL TV MAET (Ours)
MLAR Corpus Yes Yes Yes Yes Yes No
Tuning Parameters Full Full Low-Rank Low-Rank Full Ability-related
Ability Composition No No No No Yes Yes
Ability Transfer No No No No No Yes
Table 1: The difference between our proposed MAET and the methods in previous work (i.e., CPT (Hu et al., 2022), Mixture-of-Expert (MoE) (Shazeer et al., 2017), LoRA (Hu et al., 2022), Mixture-of-LoRA (MoL) (Feng et al., 2024), and Task Vector (TV) (Ilharco et al., 2023). MLAR denotes the abbreviation of multi-lingual ability-related corpus.

### v2

@@ Abstract
Multi-lingual ability transfer has become increasingly important for the broad application of large language models (LLMs). Existing work highly relies on training with the multi-lingual ability-related data, which may be not available for low-resource languages. To solve it, we propose a Multi-lingual Ability Extraction and Transfer approach, named as MAET. Our key idea is to decompose and extract language-agnostic ability-related weights from LLMs, and transfer them across different languages by simple addition and subtraction operations without training. Specially, our MAET consists of the extraction and transfer stages. In the extraction stage, we firstly locate key neurons that are highly related to specific abilities, and then employ them to extract the transferable ability-specific weights. In the transfer stage, we further select the ability-related parameter tensors, and design the merging strategy based on the linguistic and ability specific weights, to build the multi-lingual ability-enhanced LLM. To demonstrate the effectiveness of our proposed approach, we conduct extensive experiments on mathematical and scientific tasks in both high-resource and low-resource lingual scenarios. Experiment results have shown that MAET can effectively and efficiently extract and transfer the advanced abilities, outperforming competitive baseline methods and only expending 60% computational resources. Our code and data are available at https://github.com/RUCAIBox/MAET.
@@ 1 Introduction
Large language models (LLMs) have recently shown remarkable performance on various general tasks, e.g., text generation and question answering Zhao et al. (2023); OpenAI (2023); Dubey et al. (2024). Despite the success, LLMs are still struggling to solve complex tasks (e.g., mathematical reasoning), which require LLMs to possess specific advanced abilities (e.g., deductive reasoning) and knowledge (e.g., mathematical theory) Yue et al. (2024); Lu et al. (2022). To address it and further improve LLMs, existing work either collects the related data to train LLMs Du et al. (2024); Chen et al. (2024a), or merges the parameters of existing well-performed LLMs to transfer their advanced abilities into one single model Ilharco et al. (2023); Yadav et al. (2023); Yu et al. (2024a).
Despite the success, it is not easy to collect sufficient training corpus or well-trained LLMs related to specific abilities, especially in multi-lingual scenarios. Especially, some popular languages (e.g., English) have dominated the linguistic expressions of the open web data, and the amount of available domain-specific data for low-resource languages (e.g., Bengali or Telugu) is highly limited Magueresse et al. (2020); Patzelt (2024); Mirashi et al. (2024). Fortunately, existing work Zhao et al. (2024); Schäfer et al. (2024) has revealed that the learned knowledge from one language by LLMs could be inherited and leveraged by other languages. For example, Llama-series LLMs are trained mainly on English texts, while they can also solve the tasks based on other languages. Such a finding has been widely explored in either improving the overall performance of multi-lingual LLMs Schäfer et al. (2024) or enhancing fine-grained knowledge Chen et al. (2024a). However, the related work mostly relies on training with ability-related corpus in the target language, which is not always available for low-resource languages.
@@
To assess the effectiveness of our approach, we conduct the evaluation based on two comprehensive reasoning benchmarks, namely Multi-lingual Grade School Math (MGSM) Shi et al. (2023) and science tasks from multi-lingual MMLU Lai et al. (2023) as the evaluation benchmarks. According to the evaluation results, the proposed approach MAET outperforms other competitive baseline methods (e.g., continual pre-training Gururangan et al. (2020) and model merging methods with task vectors Ilharco et al. (2023), achieving the 10% relative improvement compared to the base LLM, which outperforms the competitive baseline methods. In addition, our approach can work well with relatively fewer training data, demonstrating an improved efficiency in practice.
@@ 2 Related Work
Continual Pre-training. Although LLMs still struggle in several specific tasks, e.g., complex reasoning tasks Paster et al. (2024); Shao et al. (2024) or low-resource lingual scenarios Hedderich et al. (2021); Panchbhai and Pankanti (2021). To adapt LLMs pre-trained on the general corpus to multi-lingual scenarios or specific tasks, existing work Luo et al. (2022); Taylor et al. (2022); Zhao et al. (2022); Zhang et al. (2024a) has collected the corresponding corpus to continually pre-train (CPT) LLMs. During the continual pre-training process, the mixture strategy between the general corpus and the ability-related corpus should be carefully considered to prevent hurting the general abilities of LLMs Ye et al. (2024); Xie et al. (2023); Chen et al. (2024a); Siriwardhana et al. (2024). However, previous study Chang et al. (2024); Lu et al. (2023) has pointed out that it is difficult to collect the required corpus, especially for low-resource language corpus. Therefore, synthesizing data from powerful LLMs is widely utilized to expand the task-specific training corpus Chen et al. (2021b); Yu et al. (2024b); Zhou et al. (2024a). Besides, because of the limitation of computation resources, a series of approaches Hu et al. (2022); Li and Liang (2021); Dettmers et al. (2023) only train several parameters to reduce the expenses. In this work, we focus on adapting LLMs to multilingual complex reasoning scenarios through continually pre-training LLMs on the single-lingual task-specific corpus.
@@
Knowledge Editing. According to lottery ticket hypothesis Frankle and Carbin (2019), training a small number of model parameters will achieve comparable or even better performance on downstream tasks. Existing study Du et al. (2024); Wang et al. (2024b); Gong et al. (2024) has leveraged the inner information of LLMs, to select and train the related sub-network. Besides, the probe (i.e., a newly initialized parameter) can be implemented to detect the knowledge of LLMs and process targeted repair Wang et al. (2024a); Jiang et al. (2024). However, Since calculating and selecting a sub-network of each training instance might cause the instability of the training process, several study Chen et al. (2024b); Zhang et al. (2024b) pointed out that the task-related sub-networks can be determined before the training process, and only optimizing these sub-networks can achieve better performance. In this work, we focus on editing the task-specific neurons of LLMs to improve the corresponding capacities in multi-lingual tasks.
@@
Model Merging. Given that the CPT process will bring huge computational expenses, previous work leveraged model merging techniques to integrate different abilities (e.g., mathematical reasoning and code synthesizing) into one model Yang et al. (2024); Xu et al. (2024b); Stoica et al. (2024). During the merging process, the interference between different LLMs might be conflict with each other and affect the final performance. Therefore, researchers proposed the clip Yadav et al. (2023) or randomly dropout Yu et al. (2024a) mechanism to mitigate the performance decrease. Moreover, the selection of the hyper-parameters (e.g., weight of each model) is the challenge of the model merging process, and previous work Zhou et al. (2024b); Matena and Raffel (2022) utilized the inner parameters of LLMs or external matrixes to determine the hyper-parameters. Furthermore, a series of work has studied improving the reasoning ability of LLMs in non-English scenarios by merging the reasoning-specialized model and multi-lingual model Huang et al. (2024); Yoon et al. (2024). Inspired by the above work, we try to locate the task-related sub-networks of LLMs and transfer the advanced abilities.
@@ 3 Preliminary
Despite that LLMs exhibit remarkable performance on general tasks, they still have limited advanced abilities, e.g., mathematical and scientific reasoning abilities. A typical approach to enhance these abilities is to continually pre-train (CPT) LLMs with ability-related corpus. However, such training data might not always be available or sufficient, especially for minor domains (e.g., Bengali). In this work, we focus on the task of ability extraction and transfer by continual pre-training and merging LLMs. Concretely, LLMs are trained on the collected corpus from a certain domain, and we aim to only transfer its learned advanced capabilities to target domains (Zhuang et al., 2021; Farahani et al., 2021) without further training. In this work, we study the cross-lingual scene where the linguistic-agnostic advanced ability and linguistic abilities should be extracted and transferred, to build a unified multi-lingual ability-enhanced LLM.
@@ 4 Approach
Figure 1: The framework of our approach MAET, including extraction and transfer stages. In the extraction stage, we first locate the key neurons, and utilize the single-lingual ability-related corpus and general corpus to train the LLM on these neurons to obtain the ability-related weight. Then, we remove the parameter tensors related to language knowledge in the ability weight and transfer the remaining to the base LLM. After these stages, we can obtain a powerful LLM with advanced abilities that can solve the corresponding tasks in multi-lingual scenarios.
In this section, we propose the Multi-lingual Ability Extraction and Transfer approach, named as MAET, which can effectively transfer the advanced abilities from single-lingual LLMs, to build a multi-lingual ability-enhanced LLM. The key motivation of our approach is to identify and extract ability-related neurons or weights, and transfer the target abilities into a LLM in an efficient way. The framework of MAET is presented in Figure 1.
@@ 4.1 Ability-related Weights Extraction
Locating the Key Neurons. The gradient of each neuron in LLMs can be utilized to estimate its correlation degree with specific task ability Pruthi et al. (2020); Chen et al. (2024b); Xia et al. (2024), we select those with high gradient values as key neurons. To this end, we first use the ability-related corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ to continually pre-train the LLM, while sampling a small amount to train the model can be also applied to reduce the computation consumption. During training, the LLM learns the language modeling task and each neuron is updated by the gradients associated by the training instances. Due to the high cost of calculating the accumulation of gradient at each training step, we calculate the value changes of the LLM neurons before and after the training process to approximate the importance. Formally, the importance function $I(A_{i},\theta_{j})$ of neurons can be computed as:
$I(A_{i},\theta_{j})=\sum_{d_{k}\in\mathcal{C}_{L_{0},A_{i}}^{\prime}}{\text{Grad}\left(\theta_{j},d_{k}\right)}\approx\frac{\parallel\tilde{\theta}_{j}-\theta_{j}\parallel}{\text{LearningRate}},$ (1)
where $d_{k}$ denotes the $k$-th instance of training corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ and $\tilde{\theta}_{j}$ denote the value of the $j$-th neuron of LLM after training, respectively. Based on it and inspired by previous work (Yadav et al., 2023), we rank all neurons according to their importance scores, and then select the top $k_{1}\%$ ones into the set $\mathcal{N}_{A_{i}}$ as the key neurons.
Learning Ability-related Weights. Based on the identified key neurons in $\mathcal{N}_{A_{i}}$, we further learn the ability-related parameter weights. Our motivation is to decompose the parameter weights according to their changes before and after the LLM has mastered a specific ability, which is achievable owing to the modularity and composition nature of the LLM parameter matrices (Yu et al., 2024a; Shazeer et al., 2017). First, we utilize the key neurons locating method mentioned above to extract the ability-related neuron set $\mathcal{N}_{A_{i}}$, and also obtain the language-related neuron set $\mathcal{N}_{L_{0}}$ via the same way. Then, we train the LLM with the mixture of ability-related corpus and general corpus on the key neuron set $\mathcal{N}_{A_{i}}\bigcup\mathcal{N}_{L_{0}}$ and $\mathcal{N}_{L_{0}}$ respectively, to obtain two specific models, denoted as $\text{LLM}_{A_{i},L_{0}}$ with parameters $\Theta_{A_{i},L_{0}}$ and $\text{LLM}_{L_{0}}$ with parameters $\Theta_{L_{0}}$. Next, we measure the parameter changes between the backbone and the trained models, and obtain the ability-related weights via the parameter decomposition operation as:
@@
where $\alpha$ and $\beta$ are tunable coefficients to balance the two parts of weight differences, and $\Theta_{o}$ denote the original parameters of the LLM, which serves as the reference for parameter decomposition. As we only train the parameters within the neuron set, its weight difference should preserve the knowledge about the corresponding ability. Thus, it can be regarded as the ability-related parameter representations, and is promising to transfer the ability into other LLMs by the addition operation.
@@ 4.2 Multi-lingual Ability Transfer
After obtaining the ability-related weights, we utilize them to transfer and integrate the abilities, to build a multi-lingual ability-enhanced LLM.
@@
where we use the cosine similarity to implement the similarity function $\texttt{sim}\left(\cdot\right)$. After obtaining the similarity of all tensors, we rank them in a descending order based on the similarity values, and then select the last $k_{2}\%$ parameters into the set $\mathcal{T}$ as the ability-related parameters.
@@
Approaches MLAR TPara AC AT
CPT Yes Full No No
MoE Yes Full No No
LoRA Yes Low-Rank No No
MoL Yes Low-Rank No No
TV Yes Full Yes No
MAET No Ability-related Yes Yes
Table 1: The difference between our MAET and the methods in previous work (i.e., CPT (Hu et al., 2022), Mixture-of-Expert (MoE) (Shazeer et al., 2017), LoRA (Hu et al., 2022), Mixture-of-LoRA (MoL) (Feng et al., 2024), and Task Vector (TV) (Ilharco et al., 2023). MLAR, TPara, AC, and AT denote the abbreviation of multi-lingual ability-related corpus, parameters for training, ability composition, and ability transfer.